A project specifies a concrete strength of at least 3000 psi. Materials engineers will design



the mix for a strength higher than that. Calculate the required average compressive strength of



the mix design if the standard deviation is s = 350 psi. Estimate the modulus of elasticity of the



concrete at the required average compressive strength.

1
Expert's answer
2022-03-16T07:57:01-0400

Concrete strength = 3000 psi

Standard deviation is s = 350 psi

To determine :-

The required average compressive strength of the mix design

The modulus of elasticity of the concrete at the required average compressive strength 


Step 2


The required average compressive strength of the mix design can be calculated by using the formula 

f'cr = f'c + 1.34 s

where, 

f'cr = Required compressive strength 

f'c = Specified concrete compressive strength 

s = standard deviation

Also,

Required average compressive strength for large deviation can be calculated by using the formula

f'cr = f'c + 2.33 s - 500

So, 

Larger value out of the two value is to be taken as the required compressive strength 

Modulus of Elasticity is given by formula ,

EC=57000√F'cr




Step 3


Now,

Required average compressive strength will be 

f'cr = f'c + 1.34 s

Here, 

f'c = 3000 psi

s = 350 psi

So,

f'cr = 3000 + 1.34×350

×

350

f'cr = 3000 + 469

f'cr = 3469 psi 

Also,

Required average compressive strength for large deviation will be 

f'cr = f'c + 2.33 s - 500

f'cr =3000+ 2.33×350

×

350

 - 500

f'cr = 3000 + 815.5 - 500

f'cr = 3000 + 315.5

f'cr = 3315.5 psi 

Now, 

Larger from both the value is 3469 psi 

Therefore, 

The required average compressive strength of the mix design is 3469 psi